OpenStudy (anonymous):

WILL MEDAL AND FAN!!!!!!!!!!!!!! LAST QUESTION OF MY HW!!!!!!!! Prove that a line that divides two sides of a triangle proportionally is parallel to the third side. Be sure to create and name the appropriate geometric figures.

OpenStudy (anonymous):

xD this is hard.

OpenStudy (king.void.):

Is this a part of your exams?

OpenStudy (conqueror):

. . . Why did you tell me to come here @Joel_the_boss

OpenStudy (conqueror):

Oops, I thought you were @Joel_the_boss @King.Void.

OpenStudy (anonymous):

I dont know how to start this problem

OpenStudy (king.void.):

I'm sorry but we cannot help on anything that goes on a Grade.

OpenStudy (anonymous):

This is not for a grade. But I would like help understanding the question and learning how to find the answer.

OpenStudy (conqueror):

Does this question belong to a test ? @livelovechloeee

OpenStudy (anonymous):

No. I have typed it from a study guide take home practice packet.

OpenStudy (anonymous):

 something like this?

OpenStudy (anonymous):

what grade is this?

OpenStudy (anonymous):

Im in 10th grade, this is geometry.

OpenStudy (anonymous):

xD i have not did geometry yet sorry.

OpenStudy (anonymous):

See if my drawing helps you, keep trying, and see what exactly the question is asking.

OpenStudy (anonymous):

I understand that a line is dividing the triangle in two. but what is the third side the question is referring to?

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

Given: In triangle ABC, D and E are two points, so respectively AD/DB = AE/EC. Prove: DE is parallel to BC. In triangle ABC, given, AD/DB = AE/EC. Let us assume that in triangle ABC, the point F is an intersect on the side AC. So we can see that AD/DB = AF/FC. Simplify, AE/EC = AF/FC Add 1 on both sides, (AE/EC) + 1 = (AF/FC) + 1 (AE+EC)/EC = (AF+FC)/FC AC/EC = AC/FC EC = FC From the work above, we can say that the points E and F coincide on AC, and DF coincides with DE. Since DF is parallel to BC, DE is also parallel BC.

OpenStudy (anonymous):

l